solve. Round to the nearest tenth of necessary
Gabriel walks 1/10 as fast as a train travels. It take s1.5 hr longer for him to walk 12 mi than for the train to travel 60 mi. Find the speed of each

Answer by fractalier(6550)

Let's call Gabriel's rate of walking, r. Then the rate of the train is 10r.
Let's call Gabriel's time, t. Thus the time the train takes is t - 1.5.
Then we can write (knowing that rate times time = distance):
rt = 12 and
10r(t - 1.5) = 60
A couple of ways to solve this...let's multiply out the second one and then substitute the first into the second...we get
10rt - 15r = 60
10(12) - 15r = 60
120 - 15r = 60
-15r = -60
r = 4 mph (Gabriel)
10r = 40 mph (train)